# ArtA: Automating Design Space Exploration of Spin Qubit Architectures

### v1

@@ Abstract
In the fast-paced field of quantum computing, identifying the architectural characteristics that will enable quantum processors to achieve high performance across a diverse range of quantum algorithms continues to pose a significant challenge. Given the extensive and costly nature of experimentally testing different designs, this paper introduces the first Design Space Exploration (DSE) for quantum-dot spin-qubit architectures. Utilizing the upgraded SpinQ compilation framework, this study explores a substantial design space comprising 29,312 spin-qubit-based architectures and applies an innovative optimization tool, ArtA (Artificial Architect), to speed up the design space traversal. ArtA can leverage seventeen optimization method configurations, significantly reducing exploration times by up to 99.1% compared to a traditional brute force approach while maintaining the same result quality. After a comprehensive evaluation of best-matching optimization configurations per quantum circuit, ArtA suggests universal architectural features that perform optimally across all examined circuits, emphasizing the importance of maximizing quantum gate parallelization at the expense of more crosstalk interference.
@@ I Introduction
Recognizing the intricate challenges of designing quantum processor architectures, this study initiates the first Design Space Exploration (DSE) for spin-qubit architectures, both from a current technological perspective and a future one. Therefore, in this work, we have identified a wide range of representative architectural features and abstracted them into usable input variables, resulting in $29,312$ different architectures. To facilitate this exploration, we have enhanced the compilation capabilities of SpinQ compilation framework [47] to handle all these input variables and updated the definition of the Estimated Success Probability (ESP) metric to include not only operational errors but also crosstalk and decoherence errors. These transformations and many more establish SpinQ as the first compilation and DSE framework for spin qubit architectures.
DSE processes [50, 39], which rely on a brute force approach to traverse and subsequently analyze the design space, will be impractical time-wise, especially for such large spaces. Our approach, however, employs ArtA (Artificial Architect), a built-in tool containing multiple optimization methods designed to automate the DSE process of SpinQ, thus taking less time than the brute force approach with the same quality of results. The abilities of ArtA are (a) to suggest which of the seventeen optimization configurations can find the architecture with the desired ESP the fastest compared to the brute-force approach and (b) which architectural design characteristics are key for building high-performance spin-qubit devices per quantum benchmark circuit.
Our results demonstrate ArtA’s ability to compare all optimization techniques for each quantum circuit and obtain a solution up to 99.1% faster, on average, than brute-forcing. Then, equipped with these insights, we move on to conduct a DSE analysis of $29,312$ spin qubit architectures and provide valuable insights into best design practices. Firstly, we highlight the critical need to maximize the parallelization of quantum gates rather than minimizing crosstalk between qubits. Secondly, we find that the communication method with shuttle operations is better in large-scale circuits than SWAPs, but for single-qubit gates, pulse-based methods rotations are preferred over shuttle-based. Finally, our universal architecture for all used quantum circuits suggests that the likelihood of success increases when prioritizing the parallelization of single-qubit gates instead of two-qubit gates.
@@
    The design space definition comprising of abstracted characteristics of current and, potentially, future spin qubit architectures.
@@
The remainder of this paper is structured as follows: In Section II, we discuss the scalability potential of quantum-dot spin qubits as well as the importance of addressing engineering challenges while taking into account the resulting performance of the final architecture. Then, in Section III, we motivate the need for an automated Design Space Exploration analysis for spin qubit devices and formulate four research questions for this work. After that, in Section IV, we define the design space consisting of various input architectural variables and establish the new DSE functionality of SpinQ. Another upgrade in SpinQ is the enhanced Estimated Success Probability metric, described in Section V, which is used as the Figure of Merit for our exploration. In Section VI, we introduce ArtA (Artificial Architect) and two new metrics to evaluate its performance. In the results Section VII, we first evaluate ArtA’s performance across all used quantum circuits and determine the best optimization method configuration. Then, we conduct a detailed analysis of the best architectural designs for each quantum circuit and form valuable insights for a universal architecture. We conclude our work and provide ideas for future directions in Section IX.
@@ II Challenges for scaling up spin qubit devices
Spin-qubit technologies are distinguished by their unique physical features, which position them as a highly scalable solution for quantum computing. The advantages of spin-qubits include a significantly smaller size — up to a thousand times less than other qubit technologies — combined with decades of semiconductor manufacturing expertise, long coherence times coupled with short gate durations, and high temperatures [71, 10, 30, 11, 74, 38, 63, 66, 72, 68]. At the core is the quantum dot, which contains a trapped electron(s) or hole(s) to form a physical qubit [26]. Spin qubits are manipulated electromagnetically using multiple precision-engineered gate electrodes that facilitate either single- or two-qubit operations through exact timing of pulse sequences across various quantum dot configurations. Recent studies have expanded these systems into one-dimensional and two-dimensional arrays [30, 6], exploring different structures and material combinations.
Despite the mentioned advantages, spin qubit quantum processors are not as advanced as other qubit technologies in terms of qubit counts and availability. Major technological hurdles are related to the so-called interconnect bottleneck [63], and various fabrication challenges towards scaling up [9, 5, 18, 17]. On the upside, there have been significant efforts [36, 8, 63, 31, 21, 45, 48, 65] to tackle these challenges.
However, solving them can not guarantee successful quantum algorithm executions. This is because the quality and quantity of qubits are not the only factors determining a high-performing quantum processor. The architectural constraints qubits need to comply with are equally important in how they are operated. For instance, the benefits of qubits with excellent operational fidelity can easily be outweighed by low qubit connectivity and limited natively supported quantum gates. Similarly, high-quality qubits with low crosstalk interference are not enough when they can not be addressed in parallel. These, and many more, are interlinked architectural trade-offs that affect the actual performance of the quantum processors.
@@ III Problem Statement
The need to apply design space exploration (DSE) techniques for designing and optimizing full-stack quantum computing systems, or components thereof, has been emphasized in previous work [2]. These techniques have been successfully applied across different levels of the quantum computing stack. For instance, DSE has been used to explore ion-trap quantum processors [41] and evaluate compilation techniques for superconducting qubits [54]. Such techniques can abstractly convert architectural characteristics into design variables and, in this way, quantize the design space, expressing current and future design possibilities that otherwise would be impossible or too time-consuming to physically realize. Starting from there [50, 39], the steps for a proper DSE: i) Describe the problem in terms of input variables or parameters (design choices), ii) select the performance metrics, iii) choose a global cost function as a Figure of Merit that combines different performance metrics, and iv) find models (behavioral, analytical, experimental data) that connect input parameters with metrics (or directly to the Figure of Merit). By using these DSE techniques, performance trends can be observed projecting current or future multidimensional design spaces, and points of optimality can be identified. From such an approach, it can be determined which architectural design choices (i.e., combinations of input parameter values) could provide better performance when benchmarked from small to large-scale quantum circuits.
@@
    Which optimization methods are best suited for automating the DSE? Answered in Section VII.1
@@ IV SpinQ and the design space
To answer the first two research questions, we conducted an analysis of relevant spin-qubit architectural properties with a long-term view and distilled them into numerous variables. These properties were selected to be representative of spin qubit devices as well as to create trade-offs between them. The latter will prevent maximizing one without causing a performance penalty through another (refer to Section V). The defined design space is summarized in Table 1, including their possible values. It should be noted that the DSE analysis in this work uses either discrete or categorical variables [42], as we aim to provide performance trends and rankings between optimization algorithms and architectures rather than absolute comparisons. An architecture is then characterized by one (valid) combination of variable values. Below, we describe the considered design space.
@@ IV.1 Operational gate constraints
The set of gate constraints can be vast; therefore, we decided to abstract them into the following variables that affect which (a) gate types (i.e., executed simultaneously) and (b) how many of them can be parallelized at the same time step. To highlight an example of a design trade-off, we note that the higher the degree of parallelization, the shorter the circuit depth will be, and therefore, the shorter the algorithm execution time, but it will result in higher crosstalk.
@@
    As the names suggest, these variables are boolean and signify whether specific gate types can be executed in parallel. For example, xy_z defines if X/Y can be executed simultaneously with Z gates.
@@
    This variable indicates in what manner single-qubit gates are carried out. A Global single-qubit gate implementation means the same rotation axis and angle are applied to all qubits. For qubits that do not participate in a cycle, locally applied disable instructions prevent their rotation. Then, the Semi-Global implementation is considered with the same rotation scheme as in [28] and [47], which is essentially equivalent to the implementations of the crossbar architecture proposed of [36]. In this implementation, qubit rotations are implemented semi-globally, meaning that either all qubits in odd or even column parities can be rotated at a time with the same axis and angle, and unwanted qubit rotations have to be reversed afterward by additional instructions [47]. A Local single-qubit gate implementation allows for any parallelized combination of gates at any angle to be executed in a cycle. Finally, a Sequential single-qubit gate implementation does not allow any gate parallelization, meaning that every cycle will be occupied by one gate only. It should be noted that depending on the single_qubit_impl value, only certain combinations of categories (a) and (b) are possible. For instance, in the Global and Semi-Global implementations, the xyD variable is not applicable as these implementations predetermine the upper limit of single-qubit gates allowed during a cycle.
@@ IV.2 Implementation of Z rotations:
This boolean variable also stipulates the implementation of Z rotation gates on a hardware level. The first option is a high-fidelity shuttle-based Z rotation, which is achieved with two time-sensitive qubit shuttles to and from a neighboring column [47, 28, 40, 36]. This single-qubit manipulation belongs to the so-called “hopping spins” type [67], and it is possible to achieve an arbitrary rotation axis, but in this work, we assume the aforementioned implementation proposal for simplicity. The second option is a regular pulse-based Z rotation, where Z gates are implemented similarly to X/Y gates by applying pulses. With the first choice being a unique characteristic of spin-qubits, we are interested in exploring which one of the two provides better performance across many quantum circuits. Similarly to the variables mentioned earlier, only certain combinations are possible depending on single_qubit_impl values. For example, with a Semi-Global single-qubit gate implementation and a pulse-based Z rotation gate, the zD variable is not applicable for the same reason the xyD variable is not applicable, as we explained before in category (b).
@@ IV.3 Connectivity of the device:
This variable stipulates the considered coupling graphs by addressing them through their average node degree. A coupling graph is a graph of possible direct interactions (edges) between qubits (nodes). The first coupling graph we used is a square grid (degree = 4). The second is identical to the first but adds one diagonal connection in each square (i.e., each qubit is coupled with six neighboring qubits and with three at the corners, degree = 6). Finally, the third adds the second diagonal in each square (i.e., each qubit is coupled with eight neighboring qubits and with three at the corners, degree = 8). Figure 1, presents the coupling graphs considered with color-coded edges for each of the three types. We have now reviewed all the variables related purely to hardware. Next, we will consider two variables associated also with the compilation process. More specifically, they change the routing process based on the hardware communication constraints.
Figure 1: The different coupling graphs considered in the design space. The coupling graph consists of the black lines and edges when degree = 4, black and green lines for degree = 6, and all lines when degree = 8.
@@ IV.6 Design interdependencies and verification
Figure 2: Tree of valid architectural configurations that SpinQ can handle
@@ V.2 Crosstalk term
Crosstalk has been sparsely studied for spin qubits. This is because the crosstalk effects when executing multiple gates in parallel in close proximity are still very complex to model accurately [61, 27]. Nonetheless, the presence of crosstalk is expected to have a large influence on quantum computer performance in the future, and current proposals are already suggesting techniques to mitigate it in simple forms [61]. For our purposes, crosstalk creates necessary trade-offs between architectural variables, and, as explained before, this is essential for a fair ESP calculation across architectures. For these reasons, we have devised a model to help us discover relative categorical differences between architectures, rather than accurate predictions of crosstalk effects. Our overall aim is not to approximate the actual performance of designs but to pronounce their differences by creating a performance order between them. We further assume that crosstalk should depend on local hardware characteristics between any two nearest-neighboring qubits being operated at the same time, and since this information is already included in the operational fidelity of each gate type, our crosstalk definition is a function of those. Below, we summarize our assumptions:
@@ V.3 Decoherence term
The last term of Equation 2 represents the decoherence of quantum information; hence, it creates a strong trade-off between parallelization and crosstalk effects. Since $N$ and $T_{2}$ are given, we need to define the duration time of the circuit $t$:
@@ VI ArtA (Artificial Architect)
Figure 4: Schematic overview of the initialization, optimization, and output cycle of ArtA (red box) with SpinQ (blue box).
@@
To do that, we introduce ArtA, a tool built into the SpinQ framework containing multiple optimizers. Its goal is to give answers to (a) which optimization technique performs best on the task of finding an architecture that achieves the desired ESP and (b) which architecture provides the highest ESP, both per quantum circuit.
The abstracted ArtA process is depicted in Figure 4. The initialization of ArtA starts with one or more quantum circuits, a design space such as the one described in Sec. IV, the optimization method(s) to traverse the space, and an (optional) termination condition. Inside the optimization stage, ArtA provides an architecture and circuit to SpinQ for compilation, gets the ESP value, and, based on a policy of the optimization method, selects the next architecture (in the form of input variable values) from the available “population” of architectures – a subset of the complete design space determined by the optimization algorithm itself. This process repeats until a specific termination condition (TC) (e.g., elapsed time) is met or a desired value of a specific metric is reached (in this work, ESP). Upon termination, the resulting architecture, its ESP, and ArtA metrics are stored. In the following sections, we will describe the TC and ArtA metrics.
@@
To answer the last two research questions, we will use ArtA’s (a) and (b) capabilities. In the first one, we answer the third research question by testing all the implemented optimization methods to determine which finds the highest ESP the fastest for each quantum circuit. Therefore, we will compile each circuit for all $29,312$ architectures and store their ESP. Then, we can run each method and determine the methods’ performance based on two new metrics introduced in Sec. VI.2 evaluating ArtA’s overall performance. In the second way, we will only use the best-performing optimization method but expand our circuit selection to concretely answer the last research question.

### v2

@@ Abstract
In the fast-paced field of quantum computing, identifying the architectural characteristics that will enable quantum processors to achieve high performance across a diverse range of quantum algorithms continues to pose a significant challenge. Given the extensive and costly nature of experimentally testing different designs, this paper introduces the first Design Space Exploration (DSE) for quantum-dot spin-qubit architectures. Utilizing the upgraded SpinQ compilation framework, this study explores a substantial design space comprising 29,312 spin-qubit-based architectures and applies an innovative optimization tool, ArtA (Artificial Architect), to speed up the design space traversal. ArtA can leverage 17 optimization configurations, significantly reducing exploration times by up to 99.1% compared to a traditional brute force approach while maintaining the same result quality. After a comprehensive evaluation of best-matching optimization configurations per quantum circuit, ArtA suggests specific and universal architectural features that provide optimal performance across the examined circuits. Our work demonstrates that the synergy between DSE methodologies and optimization algorithms can effectively be deployed to provide useful suggestions to quantum processor designers.
@@ I Introduction
Recognizing the intricate challenges of designing quantum processor architectures, this study initiates the first Design Space Exploration (DSE) for quantum dot spin-qubit architectures, both from a current technological perspective and a future one. Therefore, in this work, we have identified a wide range of representative architectural features and abstracted them into usable input variables, resulting in $29,312$ different architectures. To facilitate this exploration, we have enhanced the compilation capabilities of SpinQ compilation framework [51] to handle all these input variables and updated the definition of the Estimated Success Probability (ESP) metric to include not only operational errors but also crosstalk and decoherence errors. These transformations, and many more, establish SpinQ as the first compilation and DSE framework for spin qubit architectures.
Standard DSE processes [54, 43] will be impractical time-wise, especially for such large spaces, as they rely on a brute force approach to traverse and subsequently analyze the design space. Our approach, however, employs ArtA (Artificial Architect), a built-in tool containing multiple optimization methods designed to automate the DSE process of SpinQ, thus taking less time than the brute force approach with the same quality of results. The abilities of ArtA are (a) to suggest which of the seventeen optimization configurations can find the architecture with the desired ESP the fastest compared to the brute-force approach and (b) which architectural design characteristics are key for building high-performance spin-qubit devices per quantum benchmark circuit.
Our results demonstrate ArtA’s ability to compare all optimization techniques for each quantum circuit and obtain a solution up to 99.1% faster, on average, than brute-forcing. Then, equipped with these insights, we move on to conduct a DSE analysis of $29,312$ spin qubit architectures and provide valuable insights into best design practices. Firstly, we highlight the critical need to maximize the parallelization of quantum gates rather than minimizing crosstalk between qubits. Secondly, we find that the communication method with shuttle operations is better in large-scale circuits than SWAPs, but for single-qubit gates, pulse-based methods rotations are preferred over shuttle-based. Finally, our universal architecture for all used quantum circuits suggests that the likelihood of success increases when prioritizing the parallelization of single-qubit gates instead of two-qubit gates. With these, we showcase that combining DSE methodologies with optimization algorithms effectively assists in forming beneficial recommendations for quantum processor designers for current and future technological stages.
@@
    The design space definition which comprises five main characteristics of current and, potentially, future spin qubit architectures.
@@
The remainder of this paper is structured as follows: In Section II, we discuss the scalability potential of quantum-dot spin qubits as well as the importance of addressing engineering challenges while taking into account the performance of the resulting architecture. Then, in Section III, we motivate the need for an automated Design Space Exploration analysis for spin qubit devices and formulate four research questions for this work. After that, in Section IV, we define the design space consisting of various input architectural variables and establish the new DSE functionality of SpinQ. Another upgrade in SpinQ is the enhanced Estimated Success Probability metric, described in Section V, which is used as the Figure of Merit for our exploration. In Section VI, we introduce ArtA (Artificial Architect) and two new metrics to evaluate its performance. In the results Section VII, we first evaluate ArtA’s performance across all used quantum circuits and determine the best optimization method configuration. Then, we conduct a detailed analysis of the best architectural designs for each quantum circuit and form valuable insights for a universal architecture. We conclude our work and provide ideas for future directions in Section IX.
@@ II Challenges for scaling up spin qubit devices
Spin-qubit technologies are distinguished by their unique physical features, which position them as a highly scalable solution for quantum computing. The advantages of spin-qubits include a significantly smaller size — up to a thousand times less than other qubit technologies — combined with decades of semiconductor manufacturing expertise, long coherence times coupled with short gate durations, and high temperatures [76, 11, 32, 12, 79, 42, 68, 71, 77, 73]. At the core is the quantum dot, which contains a trapped electron(s) or hole(s) to form a physical qubit [28]. Spin qubits are manipulated electromagnetically using multiple precision-engineered gate electrodes that facilitate either single- or two-qubit operations through exact timing of pulse sequences across various quantum dot configurations. Studies have expanded these systems into one-dimensional and two-dimensional arrays [32, 7], exploring different structures and material combinations.
Despite the mentioned advantages, spin qubit quantum processors are not as advanced as other qubit technologies in terms of qubit counts and device availability. Major technological hurdles are related to the so-called interconnect bottleneck [68], and various fabrication challenges towards scaling up [10, 6, 19, 18]. On the upside, there have been significant efforts [38, 9, 68, 33, 23, 49, 52, 70] to tackle these challenges.
However, solving them can not guarantee successful quantum algorithm executions. This is because the quality and quantity of qubits are not the only factors determining a high-performing quantum processor. The architectural constraints qubits need to comply with in how they are operated are equally important. For instance, the benefits of qubits with excellent operational fidelity can easily be outweighed by low qubit connectivity and limited natively supported quantum gates. Similarly, high-quality qubits with low crosstalk interference are not enough when they can not be addressed in parallel. These, and many more, are interlinked architectural trade-offs that affect the actual performance of the quantum processors.
@@ III Problem Statement
The need to apply design space exploration (DSE) techniques for designing and optimizing full-stack quantum computing systems, or components thereof, has been emphasized in previous work [2, 65, 5]. These techniques or similar ones have been successfully applied across different levels of the quantum computing stack. For instance, they have been used to explore ion-trap quantum processors [45], evaluate compilation techniques [58] and qubit connectivities [41, 39, 20] for superconducting qubits. Such techniques can abstractly convert architectural characteristics into design variables and, in this way, quantize the design space, expressing current and future design possibilities that otherwise would be impossible or too time-consuming to physically realize. Starting from there [54, 43], the steps for a proper DSE: i) Describe the problem in terms of input variables or parameters (design choices), ii) select the performance metrics, iii) choose a global cost function as a Figure of Merit that combines different performance metrics, and iv) find models (behavioral, analytical, experimental data) that connect input parameters with metrics (or directly to the Figure of Merit). By using these DSE techniques, performance trends can be observed projecting current or future multidimensional design spaces, and points of optimality can be identified. From such an approach, it can be determined which architectural design choices (i.e., combinations of input parameter values) could provide better performance when benchmarked from small to large-scale quantum circuits.
@@
    Which optimization methods are best suited for automating the DSE per quantum cicruit? Answered in Section VII.1
@@ IV SpinQ and the design space
To answer the first two research questions, we conducted an analysis of relevant spin-qubit architectural properties with a long-term view and distilled them into numerous variables. These properties were selected to be representative of spin qubit devices as well as to create trade-offs between them. The latter will prevent maximizing one without causing a performance penalty through another (refer to Section V). The defined design space is summarized in Table 1, including their possible values. It should be noted that the DSE analysis in this work uses either discrete or categorical variables [46], as we aim to provide performance trends and rankings between optimization methods and architectures rather than absolute comparisons. An architecture is then characterized by one (valid) combination of variable values. Below, we describe the considered design space analytically.
@@ IV.1 Operational gate constraints
The set of gate constraints can be vast; therefore, we decided to abstract them into the following variables that affect which (a) gate types can be executed simultaneously and (b) how many of them can be parallelized at the same time step. To highlight an example of a design trade-off, we note that the higher the degree of parallelization, the shorter the circuit depth will be, and therefore, the shorter the algorithm execution time, but it will result in higher crosstalk.
@@
    As the names suggest, these variables are boolean and signify whether specific gate types can be executed in parallel. For example, xy_z defines if X/Y gates can be executed simultaneously with Z gates.
@@
    This variable indicates in what manner single-qubit gates are carried out. A Global single-qubit gate implementation means the same rotation axis and angle are applied to all qubits. For qubits that do not participate in a cycle, locally applied disable instructions prevent their rotation. Then, the Semi-Global implementation is considered with the same rotation scheme as in [30] and [51], which is essentially equivalent to the implementations of the crossbar architecture proposed of [38]. In this implementation, qubit rotations are implemented semi-globally, meaning that either all qubits in odd or even column parities can be rotated at a time with the same axis and angle, and unwanted qubit rotations have to be reversed afterward by additional instructions [51]. A Local implementation allows for any parallelized combination of gates at any angle to be executed in a cycle. Finally, a Sequential single-qubit gate implementation does not allow any gate parallelization, meaning that every cycle will be occupied by one gate only. It should be noted that depending on the single_qubit_impl value, only certain combinations of categories (a) and (b) are possible. For instance, in the Global and Semi-Global implementations, the xyD variable is not applicable as these implementations predetermine the upper limit of single-qubit gates allowed during a cycle.
@@ IV.2 Implementation of Z rotations:
This boolean variable also stipulates the implementation of Z rotation gates on a hardware level. The first option is a high-fidelity shuttle-based Z rotation, which is achieved with two time-sensitive qubit shuttles to and from a neighboring column [51, 30, 44, 38]. This single-qubit manipulation belongs to the so-called “hopping spins” type [72], and it is possible to achieve an arbitrary rotation axis, but in this work, we only assume the aforementioned variety for simplicity. The second option is a regular pulse-based Z rotation, where Z gates are implemented similarly to X/Y gates by magnetic pulse interactions. With the first choice being a unique characteristic of spin-qubits, we are interested in exploring which one of the two provides better performance across many quantum circuits. Similarly to the variables mentioned earlier, only certain combinations are possible depending on single_qubit_impl values. For example, with a Semi-Global single-qubit gate implementation and a pulse-based Z rotation gate, the zD variable is not applicable for the same reason the xyD variable is not applicable, as we explained before in category (b).
@@ IV.3 Connectivity of the device:
This variable stipulates the considered coupling graphs by addressing them through their average node degree. A coupling graph is a graph of possible direct interactions (edges) between qubits (nodes). The first coupling graph we used is a square grid (degree = 4). The second is identical to the first but adds one diagonal connection in each square (i.e., each qubit is coupled with six neighboring qubits and with three at the corners, degree = 6). Finally, the third adds the second diagonal in each square (i.e., each qubit is coupled with eight neighboring qubits and with three at the corners, degree = 8). Figure 1, presents the coupling graphs considered with color-coded edges for each of the three types. We have now reviewed all the variables related purely to hardware. Next, we will consider two variables associated also with the compilation process. More specifically, they change the routing process but based on the hardware communication constraints.
Figure 1: The different coupling graphs considered in the design space. The coupling graph consists of the black line edges when degree = 4, black and green edges when degree = 6, and all edges when degree = 8.
@@ IV.6 Design interdependencies and verification
Figure 2: Tree of valid architectural configurations that SpinQ handles.
@@ V.2 Crosstalk term
Crosstalk has been sparsely studied for spin qubits. This is because the crosstalk effects when executing multiple gates in parallel in close proximity are still very complex to model accurately [66, 29]. Nonetheless, the presence of crosstalk is expected to have a large influence on quantum computer performance in the future, and current proposals are already suggesting techniques to mitigate it in simple forms [66]. For our purposes, crosstalk creates necessary trade-offs between architectural variables, and, as explained before, this is essential for a fair ESP calculation across architectures. For these reasons, we have devised a model to help us discover relative categorical differences between architectures, rather than accurate predictions of crosstalk effects. Our overall aim is not to approximate the actual performance of designs but to pronounce their differences by creating a performance order between them. We further assume that crosstalk should depend on local hardware characteristics, and since this information is already included in the operational fidelity of each gate type, our crosstalk definition is a function of those. Below, we summarize our assumptions:
@@ V.3 Decoherence term
The last term of Equation 2 represents the decoherence of quantum information, which creates a strong trade-off between parallelization and crosstalk effects. Since $N$ and $T_{2}$ are given, we need to define the duration time of the circuit $t$:
@@ VI ArtA (Artificial Architect)
Figure 4: Schematic overview of the initialization, optimization, and output cycle when using ArtA (red box) together with the upgraded SpinQ framework (blue box).
@@
To do that, we introduce ArtA, a tool built into the SpinQ framework containing multiple optimizers. Its goal is to give answers to (a) which optimization technique performs best on the task of finding an architecture that achieves the desired ESP and (b) which architecture provides the highest ESP, both per quantum circuit and universally.
The abstracted ArtA process is depicted in Figure 4. The initialization of ArtA starts with one or more quantum circuits, a design space such as the one described in Section IV, the optimization method(s) to traverse the space, and an (optional) termination condition. Inside the optimization stage, ArtA provides an architecture and circuit to SpinQ for compilation, gets the ESP value, and, based on a policy of the optimization method, selects the next architecture (in the form of input variable values) from the available “population” of architectures – a subset of the complete design space determined by the optimization algorithm itself. This process repeats until a specific termination condition (TC) (e.g., elapsed time) is met or a desired value of a specific metric is reached (ESP in this work). Upon termination, the resulting architecture, its ESP, and ArtA metrics are stored. In the following sections, we will describe the TC and ArtA metrics in more detail.
@@
To answer the last two research questions, we will use ArtA’s (a) and (b) capabilities. In the first one, we answer the third research question by testing all the implemented optimization methods to determine which finds the highest ESP the fastest for each quantum circuit. Therefore, we will compile each circuit for all $29,312$ architectures and store their ESP. Then, we can run each method and determine the methods’ performance based on two new metrics introduced in Section VI.2 evaluating ArtA’s overall performance. In the second way, we will only use the best-performing optimization method but expand our circuit selection to concretely answer the last research question.